Summary of the August 2009 LBC Maintenance Activities

From Fernando Pedichini - 27-Aug-2009
This is a short report about our last maintenance mission from 28 to 23
August 2009:
1) Image analisys software implemented on Window PC machine (Baru Soft)
2) General maintenance of soft and hardware in all of the 5 tree-house PCs
3) Blu and Red cameras criostat full maintenance of vacuum and regeneration
with vacuum baking of external zeolith traps
4) Blu shutter enhancement trough modification to accept evaporated nitrogen
from criostat to defrost main window
5) Assembly and test of a portable vacuum turbo pump for evacuation of
cameras at the telescope.
6) Installation of new mechanics for both CCD controller of the Red channel
plus spare
7) Laboratory  test of both cameras on Blu and Red Channel using new spare
power supply
8) Delivery of a spare ccd controller power supply
9) Air venting mesh substitution on the Red channel barrel cover
10) Better positioning inside the cable chain of the RED ccd controllers
optical fibers
11) Full recabling of the power lines of the Red ccd controllers in the
cable chain at the Red Barrel with low strain cables
12) Red rotator testing with measurement of motor current NOW IT IS BELOW
1.5 Amps!
13) Both cameras remounting at prime focus stations
14) Debugging of an ethernet wire mismatch at the Blu channel power-supply
hinibiting the control of the camera
15) Full subsystems test
16) Remote desktop software istallation ontree house Windows PC for
debugging and maintenance pourposes

-- JohnHill - 17 Sep 2009
Topic revision: r2 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback